OpenStudy (bluecreep):

Elimination using multiplication heeelllp 7x+4y=-4 5x+8y=28

OpenStudy (faiithhh):

what's the question here....

OpenStudy (bluecreep):

How do I solve this by elimination using multiplication

rebeccaxhawaii (rebeccaxhawaii):

7x+4y=-4 solve for y

OpenStudy (bluecreep):

How though

TheSmartOne (thesmartone):

7x+4y=-4 5x+8y=28 Multiply the first equation by 2 so that way we can eliminate y. For example if the equation is a + b = c If we multiply both sides by 2, we get: 2(a + b) = 2(c) 2a + 2b = 2c Do that for 7x+4y=-4. Multiply it by 2 on both sides

OpenStudy (bluecreep):

I now have 14x+4y=-8

TheSmartOne (thesmartone):

A little mistake: Distributive property on the left side 2(7x + 4y) = 2(7x) + 2(4y) = 14x + 8y So the equation should be: 14x + 8y = -8 Do you understand? :)

OpenStudy (bluecreep):

Yea

TheSmartOne (thesmartone):

So let's put this new equation next to the second equation 14x + 8y = -8 5x + 8y = 28 Now, subtract the second equation from the first 14x - 5x + 8y - 8y = -8 - 28 Simplify that further :)

OpenStudy (bluecreep):

Would it be -36 or 9

OpenStudy (bluecreep):

Custom I have 9x=-20 sorry I mentioned to say -20 instead of 9

OpenStudy (bluecreep):

Because not custom

TheSmartOne (thesmartone):

-36 was correct. And so is 9 -8 - 28 That's the same thing as -1(8 + 28) = -1(36) = -36 and so now we have 9x = -36 Divide both sides by 9 to solve for x

OpenStudy (bluecreep):

X=4?

OpenStudy (bluecreep):

@TheSmartOne

TheSmartOne (thesmartone):

-36/ 4 = ? There is a negative sign. Be careful!

OpenStudy (bluecreep):

Sorry forgot to put it in

OpenStudy (bluecreep):

No how do I find y

OpenStudy (bluecreep):

Now not no

OpenStudy (bluecreep):

@TheSmartOne

TheSmartOne (thesmartone):

Plug in the value of x into any of the equations and solve for y. 7(-9) +4y=-4 Solve for y

OpenStudy (bluecreep):

Why a negative 9

OpenStudy (bluecreep):

Never mind

OpenStudy (bluecreep):

Now I have -14+4y=-4

TheSmartOne (thesmartone):

my bad it should be -4

OpenStudy (bluecreep):

So it's not what I got

TheSmartOne (thesmartone):

7(-4) +4y=-4 Solve for y

OpenStudy (bluecreep):

I now have -28+4y=-4

TheSmartOne (thesmartone):

Add 28 to both sides now :)

OpenStudy (bluecreep):

4y=24?

TheSmartOne (thesmartone):

Good, good! Now divide both sides by 4

OpenStudy (bluecreep):

Y=6?

TheSmartOne (thesmartone):

That's correct! And so we got: x = -4 y = 6 That makes the final solution in coordinate form (-4, 6) Because coordinate form is (x, y)

OpenStudy (bluecreep):

Thanks do you mind helping me with 1 more please?

TheSmartOne (thesmartone):

Sure! Since I just explained it to you, I want you to try this one on your own. Post all your work and I'll try to catch your mistakes/help you along the way :)

OpenStudy (bluecreep):

Well the problem is sorta different because it's 2x-4y=-22 3x+3y=30

TheSmartOne (thesmartone):

Ah, it's not that different. We need to multiply the first equation by a certain number, and the second equation by a different number so that way we can get the same number like 6x in both the first and second equation. So, which variable do you want to eliminate?

OpenStudy (bluecreep):

Can we eliminate 3y

TheSmartOne (thesmartone):

ok We have 4y in the first equation, and 3y in the second equation Can you find the least common multiply of 4 and 3?

OpenStudy (bluecreep):

Is it 1?

TheSmartOne (thesmartone):

List the multiples. For example, multiples of 2 are: 2, 4, 6, 8, 10, and so on

OpenStudy (bluecreep):

12 right?

OpenStudy (bluecreep):

I forgot to put the 2 in 12 sorry

TheSmartOne (thesmartone):

Correct 2x-4y=-22 3x+3y=30 That means we need to multiply the first equation by 3, and the second equation by 4

OpenStudy (bluecreep):

All of it or just the 2x-4y parts

TheSmartOne (thesmartone):

You multiply both sides and remember a(b + c) = ab + ac a(b - c) = ab - ac

OpenStudy (bluecreep):

6x-12y=66? 12+12y=120?

OpenStudy (bluecreep):

@TheSmartOne

TheSmartOne (thesmartone):

6x - 12y = -66 12x + 12y = 120 You forgot the - sign before 66, I've added it :P Not add both equations: 6x + 12x - 12y + 12y = -66 + 120 The y's should cancel, and simplify it more :)

OpenStudy (bluecreep):

BRB I have to clean

TheSmartOne (thesmartone):

Sure!